# Delving into ChatGPT usage in academic writing through excess vocabulary

### v1

@@ 1 Introduction
Recent approaches attempting to quantify the increasing use of LLMs in scientific papers fall in three groups. One group of studies employed LLM detectors, which are blackbox models trained to detect LLM writing based on ground-truth human and LLM texts (Akram, 2024; Cheng et al., 2024; Liu and Bu, 2024; Picazo-Sanchez and Ortiz-Martin, 2024). Another group of works explicitly modeled word frequency distribution in scientific corpora as a mixture distribution of texts produced by humans and by LLMs, again estimated using ground-truth human and LLM texts (Liang et al., 2024b, a; Geng and Trotta, 2024). The third group of studies relied on lists of marker words, known to be over-used by LLMs, which are typically stylistic words unrelated to the text content (Gray, 2024; Liu and Bu, 2024).
@@ 2 Results
We downloaded all PubMed abstracts until March 2024 and used all 14.2 M English-language abstracts from 2010 onwards, with only minimal filtering (see Methods). We then computed the matrix of word occurrences that shows which abstracts contain which words, resulting in a 14.2 M $\times$ 2.4 M sparse binary matrix. For each word, we obtained its occurrence frequency per year by normalizing with the total number of papers published in that year.
@@
In the following analysis, we focused on 26.6 K words with frequency $p$ above $10^{-4}$ in both 2023 and 2024. With over 1 million abstracts per year, this corresponds to ${>}100$ usages per year. Across all these words, we found many with strong excess usage in 2024 (Figure 2). Less common words with strong excess usage included delves ($r=25.2$), showcasing ($r=9.2$) and underscores ($r=9.1$), together with their grammatical inflections (Figure 2a). More common words with excess usage included potential ($\delta=0.041$), findings ($\delta=0.027$), and crucial ($\delta=0.026$) (Figure 2b).
@@
Figure 3: (a) Number of excess words per year, decomposed into the excess content words and excess style words. In each year, we show, as an example, the word with the highest frequency ratio $r$ among excess words with $p>10^{-3}$ and $r>3$. (b) Number of excess words per year, decomposed into nouns, verbs, adverbs, and adjectives.
To compare the size of excess vocabulary between years, we defined as excess words all words with $\delta>0.01$ or $\log_{10}r>\frac{\log_{10}2}{4}\log_{10}p$ where $p$ is frequency in 2024 (see dashed lines in Figure 2); these thresholds were chosen such that most words in pre-Covid years were well below (Figures S1S4). The number of excess words showed a marked rise during the Covid pandemic (up to 188 words in 2021) followed by an even larger rise (to 329) in the first quarter of 2024 (Figure 3), roughly one year after ChatGPT was released.
We manually annotated all 774 unique excess words from 2013–24 into content words, like masks or convolutional, and style words, like intricate or notably (and a small number of ambiguous words, see Methods). The excess vocabulary during the Covid pandemic consisted almost entirely of content words (such as respiratory, remdesivir, etc.), whereas the excess vocabulary in 2024 consisted almost entirely of style words (Figure 3a). Out of all 280 excess style words in 2024, 66% were verbs and 18% were adjectives. For comparison, most excess words in prior years were nouns (Figure 3b).
@@
The unprecedented increase in excess style words in 2024 allows to use them as markers of ChatGPT usage. Each frequency gap $\delta$ gives a lower bound on the fraction of abstracts that went through LLMs in 2024. For example, $\delta=0.04$ for the LLM style marker word potential means that in 2024 there were 4 percentage points more abstracts containing that word than expected based on the 2021–22 data, suggesting that at least 4% of all abstracts in 2024 went through an LLM. We reasoned that combining multiple words together can increase the lower bound. For that, we grouped together all 2024 excess style words with frequency $p<T$ and computed the frequency gap $\Delta=P-Q$ in abstracts using at least one of these words, as a function of threshold $T$ (Figure 4). Here $P$ and $Q$ are the observed and the expected frequencies of such abstracts in 2024. We obtained the highest $\Delta$ value with $T\approx 0.01$ (resulting in 209 words). The frequency gap was $\Delta_{\mathrm{rare}}=0.103$, putting the lower bound on the LLM usage in 2024 at 10% (‘rare words’ in Figure 5a). Importantly, this is only a lower bound because some of the abstracts that did go through an LLM may not contain any of the style words we used for this analysis (see Discussion).
We found that we could obtain a very similar lower bound using a non-overlapping group of only ten excess style words with high individual $\delta$ values: across, additionally, comprehensive, crucial, enhancing, exhibited, insights, notably, particularly, within. We manually selected these ten words to maximize their combined frequency gap in order to get the maximal lower bound. This group yielded a very similar frequency gap: $\Delta_{\mathrm{common}}=0.098$ (‘common words’ in Figure 5a). As the two groups were non-overlapping, this serves as an independent confirmation of the lower bound.
@@
Computational fields like computation and bioinformatics showed $\Delta=(\Delta_{\mathrm{common}}+\Delta_{\mathrm{rare}})/2\approx 0.20$ (Figure 5b). Among countries, some English-speaking countries like United Kingdom and New Zealand showed $\Delta\leq 0.03$, while countries like China, South Korea, and Taiwan showed $\Delta>0.15$ (Figure 5b). The difference between inferred genders was minor (0.08 for male and 0.07 for female, both for the first and the last authors).
Among individual journals, we found very high $\Delta$ values, e.g. 0.24 for Sensors (an open access journal published by MDPI) and 0.20 for Cureus (an open access journal with simplified review process, published by Springer Nature). We analyzed several groups of journals pooled together and found high $\Delta$ for MDPI (0.17) and Frontiers (0.17) journals. For very selective high-prestige journals like Nature, Science, and Cell, and for Nature family journals, $\Delta$ was much lower (0.06 and 0.07 respectively), suggesting that easily-detectable LLM usage was negatively correlated with perceived prestige.
To find subgroups with the strongest effect, we looked at intersections of different groups and found $\Delta=0.31$ for papers from South Korea published in Sensors and $\Delta=0.35$ for computation papers from China.
@@ 3 Discussion
Our results go beyond other studies on detecting LLM fingerprints in academic writing. Gray (2024) described a $2$-fold increase in frequency for the words intricate and meticulously in 2023, while Liang et al. (2024b) identified pivotal, intricate, showcasing, and realm as the top LLM-preferred words based on a corpus of LLM-generated text. Our study is the first to perform a systematic search for LLM marker words based on excess usage in published scientific texts. We found 280 style words with highly elevated frequencies, and indeed all the above examples appear in our list.
@@ 4.1 Dataset
We added to these data all daily PubMed (https://pubmed.ncbi.nlm.nih.gov/download) update files until the end of March 2024, and parsed the XML files in the same way, applying the same filtering criteria as González-Márquez et al. (2024). We kept the last version of the abstract for each unique PMID (PubMed ID), because the daily update files often contain updated information about the papers that were already listed in previous files. This resulted in 23 839 226 abstracts.
We then only analyzed papers with publication years from 2010 to 2024, giving us 14 182 520 abstracts for analysis.
@@ 4.2 Pre-processing
We spent substantial effort to clean the abstracts from all such contaminating strings, using over 100 regular expressions to find and eliminate them. Overall, 264 588 abstracts were affected by our cleaning procedure. We have also entirely erased 3401 abstracts of errata, corrigenda, correction, or retraction notices (identified based on titles).
We then computed a binary word occurrence matrix using CountVectorizer(binary=True) from Scikit-learn, obtaining a 14 182 520×4 119 74114182520411974114\,182\,520$\times$ 4\,119\,74114 182 520 × 4 119 741 sparse matrix. We focused the subsequent analysis on 2 388 67623886762\,388\,6762 388 676 words consisting of at least four letters and composed only out of the 26 letters of English alphabet. Note that different strings (e.g. string and strings) were treated as two distinct words.
@@ 4.4 Word annotations
We identified 774 unique excess words (surpassing thresholds on $r$ or $\delta$) from 2013 to 2024. Some of these words showed excess usage in multiple years. We sorted the list alphabetically and annotated them as content and style words while being blinded to the year in which they were selected as excess words. We assigned parts of speech (nouns, adjectives, verbs, etc.) in the same way. In case of doubt, the words were discussed between the authors. When we were not certain whether a word was content or style because of ambiguous usage, we did not label this word as either content or style.

### v2

@@ 1 Introduction
Recent approaches attempting to quantify the increasing use of LLMs in scientific papers fall in three groups. One group of studies employed LLM detectors, which are blackbox models trained to detect LLM writing based on ground-truth human and LLM texts (Akram, 2024; Cheng et al., 2024; Liu and Bu, 2024; Picazo-Sanchez and Ortiz-Martin, 2024). Another group of works explicitly modeled word frequency distribution in scientific corpora as a mixture distribution of texts produced by humans and by LLMs, again estimated using ground-truth human and LLM texts (Liang et al., 2024b, a; Geng and Trotta, 2024). The third group of studies relied on lists of marker words, known to be over-used by LLMs, which are typically stylistic words unrelated to the text content (Gray, 2024; Liu and Bu, 2024; Matsui, 2024).
@@ 2 Results
We downloaded all PubMed abstracts until the end of June 2024 and used all 14.4 M English-language abstracts from 2010 onwards, with only minimal filtering (see Methods). We then computed the matrix of word occurrences that shows which abstracts contain which words, resulting in a 14.4 M $\times$ 2.4 M sparse binary matrix. For each word, we obtained its occurrence frequency per year by normalizing with the total number of papers published in that year.
@@
In the following analysis, we focused on 26.6 K words with frequency $p$ above $10^{-4}$ in both 2023 and 2024. With over 1 million abstracts per year, this corresponds to ${>}100$ usages per year. Across all these words, we found many with strong excess usage in 2024 (Figure 2). Less common words with strong excess usage included delves ($r=28.0$), showcasing ($r=10.2$) and underscores ($r=10.9$), together with their grammatical inflections (Figure 2a). More common words with excess usage included potential ($\delta=0.045$), findings ($\delta=0.031$), and crucial ($\delta=0.029$) (Figure 2b).
@@
Figure 3: (a) Number of excess words per year, decomposed into the excess content words and excess style words. In each year, we show, as an example, the word with the highest frequency ratio $r$ among excess words with $p>10^{-3}$ and $r>3$. (b) Number of excess words per year, decomposed into nouns, verbs, and adjectives.
To compare the size of excess vocabulary between years, we defined as excess words all words with $\delta>0.01$ or $\log_{10}r>\frac{\log_{10}2}{4}\log_{10}p$ where $p$ is frequency in 2024 (see dashed lines in Figure 2); these thresholds were chosen such that most words in pre-Covid years were well below (Figures S1S4). The number of excess words showed a marked rise during the Covid pandemic (up to 189 words in 2021) followed by an even larger rise (to 382) in the first half of 2024 (Figure 3), roughly one year after ChatGPT was released.
We manually annotated all 829 unique excess words from 2013–24 into content words, like masks or convolutional, and style words, like intricate or notably (and a small number of ambiguous words, see Methods). The excess vocabulary during the Covid pandemic consisted almost entirely of content words (such as respiratory, remdesivir, etc.), whereas the excess vocabulary in 2024 consisted almost entirely of style words (Figure 3a). Out of all 319 excess style words in 2024, 66% were verbs and 16% were adjectives. For comparison, most excess words in prior years were nouns (Figure 3b).
@@
The unprecedented increase in excess style words in 2024 allows to use them as markers of ChatGPT usage. Each frequency gap $\delta$ gives a lower bound on the fraction of abstracts that went through LLMs in 2024. For example, $\delta=0.045$ for the LLM style marker word potential means that in 2024 there were 4.5 percentage points more abstracts containing that word than expected based on the 2021–22 data, suggesting that at least 4.5% of all abstracts in 2024 went through an LLM. We reasoned that combining multiple words together can increase the lower bound. For that, we grouped together all 2024 excess style words with frequency $p<T$ and computed the frequency gap $\Delta=P-Q$ in abstracts using at least one of these words, as a function of threshold $T$ (Figure 4). Here $P$ and $Q$ are the observed and the expected frequencies of such abstracts in 2024. We obtained the highest $\Delta$ value with $T\approx 0.01$ (resulting in 222 words). The frequency gap was $\Delta_{\mathrm{rare}}=0.111$, putting the lower bound on the LLM usage in 2024 at 11% (‘rare words’ in Figure 5a). Importantly, this is only a lower bound because some of the abstracts that did go through an LLM may not contain any of the style words we used for this analysis (see Discussion).
We found that we could obtain a very similar lower bound using a non-overlapping group of only ten excess style words with high individual $\delta$ values: across, additionally, comprehensive, crucial, enhancing, exhibited, insights, notably, particularly, within. We manually selected these ten words to maximize their combined frequency gap in order to get the maximal lower bound. This group yielded a very similar frequency gap: $\Delta_{\mathrm{common}}=0.110$ (‘common words’ in Figure 5a). As the two groups were non-overlapping, this serves as an independent confirmation of the lower bound.
@@
Computational fields like computation and bioinformatics showed $\Delta=(\Delta_{\mathrm{common}}+\Delta_{\mathrm{rare}})/2\approx 0.20$ (Figure 5b). Among countries, some English-speaking countries like United Kingdom and Australia showed $\Delta\leq 0.04$, while countries like China, South Korea, and Taiwan showed $\Delta>0.15$ (Figure 5b). The difference between inferred genders was minor (0.09 for male and 0.07–0.08 for female, both for the first and the last authors).
Among individual journals, we found very high $\Delta$ values, e.g. 0.24 for Sensors (an open access journal published by MDPI) and 0.19 for Cureus (an open access journal with simplified review process, published by Springer Nature). We analyzed several groups of journals pooled together and found high $\Delta$ for MDPI (0.19) and Frontiers (0.18) journals. For very selective high-prestige journals like Nature, Science, and Cell, and for Nature family journals, $\Delta$ was much lower (0.06 and 0.08 respectively), suggesting that easily-detectable LLM usage was negatively correlated with perceived prestige.
To find subgroups with the strongest effect, we looked at intersections of different groups and found $\Delta=0.33$ for papers from South Korea published in Sensors and $\Delta=0.38$ for computation papers from China.
@@ 3 Discussion
Our results go beyond other studies on detecting LLM fingerprints in academic writing. Gray (2024) described a $2$-fold increase in frequency for the words intricate and meticulously in 2023, while Liang et al. (2024b) identified pivotal, intricate, showcasing, and realm as the top LLM-preferred words based on a corpus of LLM-generated text. Our study is the first to perform a systematic search for LLM marker words based on excess usage in published scientific texts. We found 319 style words with highly elevated frequencies, and indeed all the above examples appear in our list.
@@ 4.1 Dataset
We added to these data all daily PubMed (https://pubmed.ncbi.nlm.nih.gov/download) update files until the end of June 2024, and parsed the XML files in the same way, applying the same filtering criteria as González-Márquez et al. (2024). We kept the last version of the abstract for each unique PMID (PubMed ID), because the daily update files often contain updated information about the papers that were already listed in previous files. This resulted in 24 105 518 abstracts.
We then only analyzed papers with publication years from 2010 to 2024, giving us 14 448 711 abstracts for analysis.
@@ 4.2 Pre-processing
We spent substantial effort to clean the abstracts from all such contaminating strings, using over 100 regular expressions to find and eliminate them. Overall, 270 189 abstracts were affected by our cleaning procedure. We have also entirely erased 3431 abstracts of errata, corrigenda, correction, or retraction notices (identified based on titles).
We then computed a binary word occurrence matrix using CountVectorizer(binary=True) from Scikit-learn, obtaining a 14 448 711×4 179 57114448711417957114\,448\,711$\times$ 4\,179\,57114 448 711 × 4 179 571 sparse matrix. We focused the subsequent analysis on 2 417 35424173542\,417\,3542 417 354 words consisting of at least four letters and composed only out of the 26 letters of English alphabet. Note that different strings (e.g. string and strings) were treated as two distinct words.
@@ 4.4 Word annotations
We identified 829 unique excess words (surpassing thresholds on $r$ or $\delta$) from 2013 to 2024. Some of these words showed excess usage in multiple years. We sorted the list alphabetically and annotated them as content and style words while being blinded to the year in which they were selected as excess words. We assigned parts of speech (nouns, adjectives, verbs, etc.) in the same way. In case of doubt, the words were discussed between the authors. When we were not certain whether a word was content or style because of ambiguous usage, we did not label this word as either content or style.